wiki:UC20181001

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 10/01/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), Misty Bentz (GSU), John Bally (CU), Chip Kobulnicky (UWy), Eric Bellm (UW), Mary Beth Kaiser (JHU), John Wisniewski (OU), Ben Williams (UW), Kal Kadlec (UW), Anne Verbiscer (UVa), John Wilson (UVa), Sarah Tuttle (UW), Joanne Hughes (Seattle U)


User feedback and comments from institutional representatives:

  • UVa - nothing to report
  • JHU - nothing to report
  • Georgia State - nothing to report
  • NAPG - no report
  • NMSU - users are interested in learning more about status of DIS (i.e. expectation that the vacuum will continue to deteriorate and require frequent servicing)- this was discussed in more detail below
  • Wyoming - nothing to report
  • Colorado - nothing to report
  • Oklahoma - nothing to report
  • Washington - same concerns about DIS as NMSU
  • Seattle - mirror covers were having an issue but it got resolved

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 9/1/18 through 9/27/18

0) Overview

We completed summer shutdown and resumed science observations minus DIS. On-sky engineering went well and the few problems we had were corrected quickly.

1) Telescope

  • M1 washed
  • M2 recoated w/ protected Aluminum
  • M3 recoated w/ protected Aluminum (still needs re-polishing in the future due to acid etching that occurred previously during stripping)
  • M2 received actuator servicing and new motors plus electronic ground mod
  • M2 and M3 limit switches upgraded

2) Instruments

DIS: Red camera new feedthrough installed, leak checked and on vacuum pump. DIS blue is on vacuum pump. We anticipate vacuum to deteriorate again and will require frequent servicing or extended maintenance this year. Summer shutdown work included new o-rings, ion pumps, cryotigers, collimation check.

TripleSpec is operational.

Agile is operational.

Echelle is operational but IOL degraded from last report. Vacuum serviced over summer.

NIC-FPS is operational.

ARCTIC is operational, improvements made to filter wheel. Vacuum serviced over summer with new ION pump.

3) ARCSAT

In use for science, guider scope was realigned, new DFM software installed.


Additional telescope and instrument discussion:

Overall we are pretty happy with the state of the mirrors post-shutdown. Because M3 not was repolished it is not as good as had been originally hoped, but it still appears to be better than what it was, and not degrading as quickly as it did with the bare aluminum coating. There was a mirror cover issue shortly after shutdown but it got resolved. The telescope is currently behaving and the instruments are mostly running well.

Some work was done on the ARCTIC filter wheel (problem trying to fix: repeatability w/filter positioning). We're still trying to understand what we're seeing after the recent software and mechanical changes. Prior to these fixes the wheel would drift after many moves; we believe that issue is now resolved. There was also an issue when coming out of shutdown where multiple homes were needed to correct the filter wheel homing routine; that also has been corrected.

ARCSAT has been back up for science for a while now. The new guider scope/camera are now installed but not yet aligned, but they are working. Its position still needs to be adjusted to reduce vignetting so the finder scope will be relocated to a new location this week or next. The software is showing its usual bugs but nothing too significant or worrisome (and they may not be easily fixable given that we are using commercial software).

The DIS cameras are currently on the vacuum pumps - red is not as good as we would like after pumping over the weekend and all last week. Steps have been taken to further outgas any water vapor in the dewars since the instrument had been open for a long time (2 weeks) while repairs to the electrical feedthroughs were going on. The current best guess is that the instrument will go back together tomorrow (10/02/18), it will be on the vacuum pump for one more day on observing level, and then it will be aligned, focused, and returned to science Wednesday evening (10/03/18) at the earliest. Ben reminded us that there is an on-site observer scheduled to use DIS pretty soon so the sooner we can be notified of any additional delays to the schedule, the better. Even though the major vacuum leak has been fixed on the red camera, whatever slow leak has caused the recurring contamination to appear on the field lenses has not been addressed. We expect that this problem will continue to occur and that we will have to do monthly (?) pumpings to mitigate it. A major redesign of the dewars would be needed to completely eliminate it.


Q4 3.5-m scheduling:

We got through September without DIS thanks to Russet’s heroic efforts and some last minute observers stepping up. DIS will be back on the telescope on Wednesday 10/03 (hopefully) and the next DIS users are scheduled for that same night. Some payback time was already assigned to those who were displaced due to the unavailability of DIS and there is still some open time available. Nancy encouraged the UC reps to encourage their users, particularly those who missed out on earlier DIS observations, to request some of this time if they can make use of it.


Q4 ARCSAT scheduling:

The Q4 ARCSAT schedule has been published.


NA2 Baffle Project

John Wilson reported that there was a fair amount of progress on this over the past month. When John was on site at APO Mark helped him visually inspect the NA2 baffle from the vantage point of the primary, which was helpful to see the outer diameter of the baffle tube. It does match the one drawing they have of the spacings. The full set of drawings of the fork bearings was located; this was the last piece needed for the completion of the SolidWorks model. The UVa team has made progress in Zemax and now they have a working model of a proxy of an expanded FOV NA2 baffle, which allows the light that goes to the guider to go to one side of the square extruded baffle. They will next explore the vignetting of the guider field (John will iterate with Bill and Russet on the maximum allowable vignetting on the guider) and the trades between having an open slit on one side vs. stray light induced on science focal plane.


KOSMOS update

The UW team has powered up camera but is having trouble getting the torrent (camera controller) to read; they are talking with the NOAO folks about this and expect this issue to be resolved shortly. They want to put the dewar on the vacuum but they want to make sure that the temperature readout is correct before cooling it down. They have a design for a slit viewer camera in hand but recommend that we have a Preliminary Design Review (PDR) first before purchasing hardware. Mark invited the UC reps (or any other interested users) to sit in on the PDR if desired. We want to make sure that as we move forward the instrument meets as many of the users' needs as possible.


TEWG Update

There was not much activity to report on from the Training and Education Working Group this month. Nancy hopes that the white paper will be completed by the end of CY18.


Response to Call for White Papers for 2020 Decadal Survey

Nancy contacted several observatory directors regarding a possible white paper describing the roles that 3-4m class telescopes can play in observational astronomy in the next decade. She has received some positive responses so at this point plans to proceed with drafting something that people can provide input and comments on. She plans to get something to the UC reps in 2 weeks so that we can discuss it on the next UC telecon in 4 weeks.


ARC Board of Governors Meeting

The ARC BoG meeting is on Monday November 5. As has been done previously, we would like to showcase science highlights from the past year from all ARC partner and leasing institutions. Nancy will likely not have time to verbally explain each highlight during her presentation, but she would like to have a comprehensive packet that she can provide to the Board members for their reference. It is extremely important for the Board members to see that their investment in the 3.5m continues to benefit their institutions. To that end, Nancy is requesting that these science highlights be sent to her no later than October 15. The specifics are as follows:

  • deadline: October 15
  • format: Powerpoint, Keynote, or PDF
  • length: one slide
  • content:
    • name(s) and institution of researcher(s)
    • title of project
    • one or more graphic showing the result (image, spectrum)
    • a text block explaining the result
    • paper citation if there is one

Individuals can send their slides to Nancy directly, or an institution can designate someone (e.g. the Users Committee rep!) to collect the slides and send them to Nancy as a single packet.


ACTION ITEMS:

Action items from previous meetings:

  • Bill and/or Russet will update the Instrument Status web page to reflect the current status of the ARCTIC filter wheel positioning (need to rehome twice). STATUS: CLOSED. This was done. Since that time the problem was corrected, and the page was again updated to reflect the current status.
  • Russet will update the calibration section of the ARCTIC users manual once the Q3/Q4 scheduling tasks are completed. STATUS: OPEN.
  • All member institutions and leasing partners should collect science highlight slides and send them to Nancy no later than October 15. STATUS: OPEN.

New action items from this meeting:

  • UC reps should remind their users who were displaced by DIS unavailability that there is some OPEN time remaining in the Q4 schedule that they can ask for.
  • Nancy will send out a draft of a white paper for the 2020 Decadal for UC feedback prior to the next telecon.

Next meeting:

The next Users Committee telecon will be on Monday October 29 at 10:30 am MDT.

Note: See TracWiki for help on using the wiki.